What is the area of triangle with vertices


(1,-1,2), (2,1,-1), (3,-1,2)

Grade:11

5 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Deeksha

First calculate the sides of the triangle

let A=(1,-1,2), B=(2,1,-1), C=(3,-1,2)

AB=(2-1)i + (1-(-1))j + (-1-2)k

AB=i + 2j - 3k

Similarly

side AC= 2i

side BC=i - 2j + 3k

area of triangle= 0.5 |(cross product of AB and AC)|

                    = 0.5|ABxAC|

                   =0.5|-4k-6j|

                   =0.5(√42 + 62)

                   =√15

                 = 3.87

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

let A,B,C are the vertices of triangle

A,B,C are (1,-1,2) , (2,1,-1) , (3,-1,2) respectively

position vector of A is i-j+2k,

position vector of B is 2i+j-k & 

position vector of C is 3i-j+2k

now ,

             AB =B-A=i+2j-3k  &  BC=B-C=i-2j+3k

area of triangle is (AB)*(BC)= -6j -4k              (* represents cross product)

                                      =sqrt(36+16)      in magnitude

                                      =2sqrt13 square units

rajan jha
13 years ago

THE AREA OF THE TRIANGLE WILL BE 1/2(DETERMINANTOF THE COOORDINATES);

LIKE MY ANSWER?

Devasish Bindani
13 years ago

let A=(1,-1,2)

    B=(2,1,-1)

    C=(3,-1,2)

by distance formula AB=(14)1/2  BC=(14)1/2  CA=2

height=(15)1/2

we know that area of triangle is 1/2*(base)*(height)

thus area ABC=1/2*2*(15)1/2 =(15)1/2

FITJEE
10 years ago

3.87